Gaseous hydrogen has a density of 0.0899 g/L at 0 °C and at 1 atm.  How many liters would you need if you wanted 1.0078 g of hydrogen? 


1
Expert's answer
2021-05-31T02:01:16-0400

Molar Mass of H = 1.00784

D = M/V

0.0899 = 1.00784/V

V = 1.00784/0.0899

= 11.211

M1/V1 = M2/V2

0.0899/11.211 = 1.0078/V2

0.00802 = 1.0078/V2

= 125.661L